Consider the following IVP

y + 5y + y  =   f (t),     y(0)  =  3,    y(0)  =  0,

where

f (t)  =
{ 8 0  ≤  t  ≤  2π
cos(7t) t  >  2π

(a) Find the Laplace transform F(s)  =  ℒ { f (t)} of  f (t).
(b) Find the Laplace transform Y(s)  =  ℒ {y(t)} of the solution y(t) of the above IVP.Consider the following IVP y + 5y + y = f(t), y(0) = 3, y(0) = 0, where f(t) {85(70) > 2.5 Osts 20 cos(7t) t > 2 L {f(t)}
Answer #1

Given IvP is y+5y ty=flt) Y(0) = 3, yllo) = 0 and f(t) = 8,osts21 Coo 7t, t>271 { f(t) can be written as $4+) = 8 (u(to) -:h[uct)] = eos = 1 L[ultan)] = 2 L[657t] =5 diyo 52449 L[coolt a(4-27) e2TS s 52+ug [[+1)] = 8(1)-9 1820s)+ e 2ns Ehtes L[f((52 Ys)-35-1)+5 (sYO-3) + (5) = 861-829s) teins (5 149) = (52 +55+1) kis) - 35-15= 861-62ns) tens (vo) =(52+55+1) YCS) = 23+3Please let me know if you have any issues or queries with the solution